How many slits per centimeter does a grating have if the third-order occurs at an 19.0? angle for 530-nm light?

Express your answer using three significant figures
Answer #1

d*sin(theta) = m*lambda

d*sin(19) = 3*530*10^-9

d = 4.8834*10^-6 m = 4.8834*10^-4 cm

slits per cm = 1/d = 1/4.8834*10^-4 = 2.05*10^3 lines/cm